OpenStudy (anonymous):

Find dy/dt by implicit differentiation and evaluate the derivative at the point (1,1) x^3 + y^3= 2xy

OpenStudy (anonymous):

I kind of started it but I got lost really quickly

OpenStudy (anonymous):

is the question is finding dy/dx or dy/dt ?? Please check it again

OpenStudy (anonymous):

oh Im sorry! dy/dx

OpenStudy (phi):

can you do d/dx x^3 ?

OpenStudy (anonymous):

3x^2?

OpenStudy (anonymous):

I got 3x^2+ 3y^2 * dy/d

OpenStudy (phi):

yes now d/dx y^3 it is almost the same 3y^2 dy/dx (the dy/dx must go in there)

OpenStudy (anonymous):

dy/dx* I just dont know what the 3xy turns into

OpenStudy (phi):

for x*y use the product rule

OpenStudy (phi):

d(xy) = x * d/dx y + y * d/dx x

OpenStudy (anonymous):

what is the first and what is the second function? is he first 2x and the second y?

OpenStudy (phi):

the order will not matter

OpenStudy (anonymous):

In case of dy/dx, you can solve by implicit differentiation as following \[d(x^3 + y^3) = d(2xy)\] \[3x^2 + 3y^2.y \prime = 2 (y + xy \prime)\] \[y \prime (3y^2-2x) = 2y - 3x^2 \rightarrow y \prime = \frac{ 2y - 3x^2 }{ 3y^2 - 2x }\] Evaluate at (1,1) \[y \prime = \frac{ 2 - 3 }{ 3 - 2 } = -1\]

OpenStudy (phi):

and I would leave the 2 out front (it will multiply the answer) 2 * d(xy)

OpenStudy (anonymous):

So is the derivative of 2xy 2x * y' + 2y? If it isnt, I have no clue then

OpenStudy (phi):

that looks good

OpenStudy (anonymous):

So then does it become dy/dx=-3x^2-3y^2 - 2x

OpenStudy (phi):

in detail: \[ 2 \frac{d}{dx} (xy) = 2\left(x\frac{d}{dx} y + y \frac{d}{dx}x\right) \\ 2 x \frac{dy}{dx} + 2y \frac{dx}{dx} \\ 2 x \frac{dy}{dx} + 2y \]

OpenStudy (anonymous):

Idk how to clean it up and simplify it

OpenStudy (phi):

you sub in (1,1)

OpenStudy (phi):

you should (at the start) have \[ 3 x^2 + 3y^2\frac{dy}{dx} = 2x \frac{dy}{dx} + 2y \] I would plug in x=1, y=1 and simplify to find dy/dx

OpenStudy (anonymous):

like how do you find out what dy/dx is because that is the first part of the problem? I have dy/dx on both sides of the equation

OpenStudy (anonymous):

isnt is supposed to be dy/dx= something?

OpenStudy (phi):

btw *** So then does it become dy/dx=-3x^2-3y^2 - 2x *** I don't know how you got this. ** isnt is supposed to be dy/dx= something? *** Yes. you treat dy/dx just like an unknown variable (e.g. "x") that you "solve for" But it might be faster to first put in 1 for x and y first...

OpenStudy (anonymous):

okay so then I get 6 dy/dx= 2dy/dt+ 2?

OpenStudy (phi):

Do you understand how we got to \[ 3 x^2 + 3y^2\frac{dy}{dx} = 2x \frac{dy}{dx} + 2y \] ? now sub in (1,1) (which means x=1, y=1) \[ 3 + 3 \frac{dy}{dx}= 2 \frac{dy}{dx} + 2\] if it helps, think of this as \[ 3 + 3x = 2x + 2 \] solve for x, and rename x as \( \frac{dy}{dx} \)

OpenStudy (phi):

***okay so then I get 6 dy/dx= 2dy/dt+ 2?*** your algebra is not correct.

OpenStudy (anonymous):

how?

OpenStudy (phi):

3+3x=2x+2 can you solve this for x ?

OpenStudy (phi):

add -3 to both sides, then add -2x to both sides then simplify

OpenStudy (anonymous):

x=-1

OpenStudy (phi):

now all I did was temporarily rename \(\frac{dy}{dx}\) as x to make the equation look simpler. in other words, you would do the same thing with \[ 3+3\frac{dy}{dx}=2\frac{dy}{dx}+2 \]

OpenStudy (phi):

add -3 to both sides, add -2 \(\frac{dy}{dx}\) to both sides

OpenStudy (anonymous):

So when you find the derivative at (1,1) the answer is dy/dx= -1?

OpenStudy (phi):

yes, the answer is \[ \frac{dy}{dx}= -1 \]at the point (1,1)

OpenStudy (anonymous):

So there isnt two different answers for this problem?

OpenStudy (anonymous):

the differentiation and the derivative at the point is all one answer?

OpenStudy (phi):

If you scroll up and look at linh's post, he shows how to solve for dy/dx (he used y' for dy/dx) in terms of x and y, then subbed in x=1 and y=1 at the end.

OpenStudy (phi):

** the differentiation and the derivative at the point is all one answer? *** the general answer for dy/dx is in terms of x and y see linh's post. once you replace x and y with 1, you get dy/dx at that point... it will be a number. (-1 in this case)

OpenStudy (anonymous):

oh okay thank you

OpenStudy (phi):

can you do the algebra to solve for dy/dx in terms of x and y, starting with \[3 x^2 + 3y^2\frac{dy}{dx} = 2x \frac{dy}{dx} + 2y \] it looks a bit more complicated... But reading the question, I think they want to first solve for dy/dx, and then substitute in (1,1) (I skipped that step, and substituted first, because I was solving for dy/dx at (1,1) )